# An Analysis of Causal Effect Estimation using Outcome Invariant Data Augmentation

### v1

@@ 1 Introduction
A classical problem in machine learning is that of regression—using i.i.d. samples from some fixed, unknown distribution $\mathbb{P}_{{X,Y}}$, we predict outcome $Y$ values for unlabelled treatment $X$ values. The use of regularization techniques is crucial for this task to achieve good generalization from training to test data [vapnik]. Data augmentation (DA) [da-reg, invariance_benefit] is one such method, where each sample is randomly perturbed multiple times to grow the dataset size. However, these regression models cannot generally be interpreted causally as the statistical relationship between $X$ and $Y$ may arise from shared common causes, known as confounders, rather than from $X$ influencing $Y$. Removing such confounders requires independently assigning values of $X$ during data generation, known as an intervention [eoci, Pearl_2009].
Unfortunately, we seldom have access to the data generation process to be able to intervene on variables. A common workaround is to use auxiliary variables to correct for unobserved confounders [backdoor, specIV, prox-causal-learning]. One such approach is that of instrumental variables (IVs) that represent certain conditional independences in the system which can be used to identify the causal effect of $X$ on $Y$ [iv_arthur, iv_rui, iv_niki]. Alas, IVs too are generally hard to find in may popular applications such as computer vision and natural language processing, motivating the need for more accessible ways to mitigate unobserved confounding.
@@ 2.1 Intervention for causal effect estimation
Minimizing Eq.˜3 is meaningful in two important cases where ERM fails: (i) Causal effect estimation: The minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ of Eq.˜3 gives us an unbiased estimate of the average treatment effect (ATE) [backdoor] $\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)}\left[Y\;\middle|\;X={\mathbf{x}}\right]=f\left({\mathbf{x}}\right)$ that measures the causal influence of $X$ on $Y$. (ii) Robust prediction: ATE based prediction of $Y$ values for unlabelled $X$ values is robust in the sense that it can generalize across arbitrary OOD treatment interventions or shifts in the treatment distribution [causal-ood-generalizatoin]. Consequently, the causal risk minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ is also a robust predictor over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$. Specifically, $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ minimizes the worst-case ERM objective over the set $\mathcal{P}$} of all possible intervention distributions $\mathbb{P}_{{X^{\prime}}}$ over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$ [causal-ood-generalizatoin], i.e. for ${\mathcal{P}}\coloneqq\left\{\mathbb{P}_{{X^{\prime}}}\;\middle|\;\operatorname{supp}\left(\mathbb{P}_{{X^{\prime}}}\right)\subseteq\operatorname{supp}\left(\mathbb{P}_{{X}}^{\mathfrak{M}}\right)\right\}$,
@@ 3 Faithfulness and Outcome Relevance in IVs
Despite problem miss-specification for a IVL $Z$, the target function $f$ remains a minimizer for the IV risk in Eq.˜5. Albeit, potentially not unique—for example, a linear $h$ with squared loss leads to an under-determined problem in Eq.˜5. We therefore propose the following regularized version of the IV risk for such an IVL setting,
@@
where $\alpha>0$ is the regularization parameter. The ERM risk as a penalty allows our estimations to have good predictive performance while the IV risk encourages solution search within the subspace where we know $f$ to be present. We refer to minimising the risk in Eq.˜7 as IVL regression.
@@ 4 Causal Effect Estimation using Data Augmentation
Given samples for only $\left(X,Y\right)$ and some valid DA parameterised by $G$, the task is to improve predictive performance across interventions and reduce confounding bias in ATE estimates. We now make two observations in the following sections and state the respective results that follow thereof.
@@ 4.1 Data augmentation as a soft intervention
That is, DA strictly reduces confounding bias in ATE estimate iff the induced intervention perturbes $X$ along spurious features. Importantly, Theorem˜3 suggests that lower confounding bias is not a ‘free lunch’ with outcome invariance of DA and practitioners may need domain knowledge to construct DA that targets spurious features. Fortunately however, Theorem˜3 also suggests that with outcome invariance, DA should not perform worse than ERM. We say that DA+ERM dominates ERM on causal estimation [domination, p. 48]. Practitioners may therefore be advised to generously use such DA, as it achieves regularization in the worst case, and mitigates confounding bias as a ‘bonus’ in the best case.
@@ 6 Experiments
In addition to standard ERM, DA and IV regression, our baselines include DRO [dro], invariant risk minimization (IRM) [arjovsky2020invariant], invariant causal prediction (ICP) [icp], regularization with invariance on causal essential set (RICE) [rice], variance risk extrapolation (V-REx) and minimax risk extrapolation (MM-REx) [riskextrap]. We also include the causal regularization method by Kania and Wit [causal-regularization2] and the $\ell_{1},\ell_{2}$ approaches by Janzing [causal-regularization]. We discretise $G$ if the method accepts only discrete variables. For IVL regression, we select the regularization parameter $\alpha$ in a variety of ways, including vanilla cross validation (CV), level-based CV (LCV) and confounder correction (CC) as described in LABEL:app:implementation-details. Other implementation details are provided in LABEL:app:experiments, and the code to reproduce our results is publicly released at https://github.com/uzairakbar/causal-data-augmentation.

### v2

@@ 1 Introduction
A classical problem in machine learning is that of regression—using i.i.d. samples from some fixed, unknown distribution $\mathbb{P}_{{X,Y}}$, we predict outcome $Y$ values for unlabeled treatment $X$ values. The use of regularization techniques is crucial for this task to achieve good generalization from training to test data [vapnik]. Data augmentation (DA) [da-reg, invariance_benefit] is one such method, where each sample is randomly perturbed multiple times to grow the dataset size. However, these regression models cannot generally be interpreted causally as the statistical relationship between $X$ and $Y$ may arise from shared common causes, known as confounders, rather than from $X$ influencing $Y$. Removing such confounders requires independently assigning values of $X$ during data generation, known as an intervention [eoci, Pearl_2009].
Unfortunately, we seldom have access to the data generation process to be able to intervene on variables. A common workaround is to use auxiliary variables to correct for unobserved confounders [backdoor, specIV, prox-causal-learning]. One such approach is that of instrumental variables (IVs) that represent certain conditional independences in the system which can be used to identify the causal effect of $X$ on $Y$ [iv_arthur, iv_rui, iv_niki]. Alas, IVs too are generally hard to find in many popular applications such as computer vision and natural language processing, motivating the need for more accessible ways to mitigate unobserved confounding.
@@ 2.1 Intervention for causal effect estimation
Minimizing Eq.˜3 is meaningful in two important cases where ERM fails: (i) Causal effect estimation: The minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ of Eq.˜3 gives us an unbiased estimate of the average treatment effect (ATE) [backdoor] $\mathbb{E}^{\mathfrak{M};\operatorname{do}\left(X\coloneqq{\mathbf{x}}\right)}\left[Y\;\middle|\;X={\mathbf{x}}\right]=f\left({\mathbf{x}}\right)$ that measures the causal influence of $X$ on $Y$. (ii) Robust prediction: ATE based prediction of $Y$ values for unlabeled $X$ values is robust in the sense that it can generalize across arbitrary OOD treatment interventions or shifts in the treatment distribution [causal-ood-generalizatoin]. Consequently, the causal risk minimizer $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ is also a robust predictor over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$. Specifically, $\hat{h}^{\mathfrak{M}}_{\text{CR}}$ minimizes the worst-case ERM objective over the set $\mathcal{P}$} of all possible intervention distributions $\mathbb{P}_{{X^{\prime}}}$ over the support of $\mathbb{P}_{{X}}^{\mathfrak{M}}$ [causal-ood-generalizatoin], i.e. for ${\mathcal{P}}\coloneqq\left\{\mathbb{P}_{{X^{\prime}}}\;\middle|\;\operatorname{supp}\left(\mathbb{P}_{{X^{\prime}}}\right)\subseteq\operatorname{supp}\left(\mathbb{P}_{{X}}^{\mathfrak{M}}\right)\right\}$,
@@ 3 Faithfulness and Outcome Relevance in IVs
Despite problem misspecification for a IVL $Z$, the target function $f$ remains a minimizer for the IV risk in Eq.˜5. Albeit, potentially not unique—for example, a linear $h$ with squared loss leads to an under-determined problem in Eq.˜5. We therefore propose the following regularized version of the IV risk for such an IVL setting,
@@
where $\alpha>0$ is the regularization parameter. The ERM risk as a penalty allows our estimations to have good predictive performance while the IV risk encourages solution search within the subspace where we know $f$ to be present. We refer to minimizing the risk in Eq.˜7 as IVL regression.
@@ 4 Causal Effect Estimation using Data Augmentation
Given samples for only $\left(X,Y\right)$ and some valid DA parameterized by $G$, the task is to improve predictive performance across interventions and reduce confounding bias in ATE estimates. We now make two observations in the following sections and state the respective results that follow thereof.
@@ 4.1 Data augmentation as a soft intervention
That is, DA strictly reduces confounding bias in ATE estimate iff the induced intervention perturbs $X$ along spurious features. Importantly, Theorem˜3 suggests that lower confounding bias is not a ‘free lunch’ with outcome invariance of DA and practitioners may need domain knowledge to construct DA that targets spurious features. Fortunately however, Theorem˜3 also suggests that with outcome invariance, DA should not perform worse than ERM. We say that DA+ERM dominates ERM on causal estimation [domination, p. 48]. Practitioners may therefore be advised to generously use such DA, as it achieves regularization in the worst case, and mitigates confounding bias as a ‘bonus’ in the best case.
@@ 6 Experiments
In addition to standard ERM, DA and IV regression, our baselines include DRO [dro], invariant risk minimization (IRM) [arjovsky2020invariant], invariant causal prediction (ICP) [icp], regularization with invariance on causal essential set (RICE) [rice], variance risk extrapolation (V-REx) and minimax risk extrapolation (MM-REx) [riskextrap]. We also include the causal regularization method by Kania and Wit [causal-regularization2] and the $\ell_{1},\ell_{2}$ approaches by Janzing [causal-regularization]. We discretize $G$ if the method accepts only discrete variables. For IVL regression, we select the regularization parameter $\alpha$ in a variety of ways, including vanilla cross validation (CV), level-based CV (LCV) and confounder correction (CC) as described in LABEL:app:implementation-details. Other implementation details are provided in LABEL:app:experiments, and the code to reproduce our results is publicly released at https://github.com/uzairakbar/causal-data-augmentation.